OpenStudy (anonymous):

find theadditive inverse of 5 on the 12- hour clock

OpenStudy (perl):

on the 12 hour clock you only have 1,2,3,4,5,6,7,8,9,11 and 12 (or zero)

OpenStudy (perl):

5 + ? = 12

OpenStudy (perl):

that number will be the additive inverse of 5 in general to solve a + x = 0 , x is the additive inverse of a

OpenStudy (solomonzelman):

or, just mathematically, -5 would be the answer, not that the time can go backwards, so it's not.

OpenStudy (kainui):

I think both of you sort of put the pieces out there well, but to summarize, the additive inverse is clearly -5 normally.\[5-5=12\] but on a clock you can always add 12 to any number and get the same number back because it's cyclic, so we take -5 and add 12 to it \[-5+12 = 7\] And that's the more common way of representing.